OpenStudy (anonymous):

A tree is supported a wire anchored in the ground 11 feet from its base. The wire is 2 feet longer than the height that it reaches on the tree. Find the length on the wire.

OpenStudy (wolf1728):

Here is a diagram

OpenStudy (wolf1728):

Let's call the tree height x the length of the wire (the hypotenuse) w and the distance from the tree 'd' which is eleven feet. So, going by the Pythagorean Theorem w² = 11² + x² we are told that w is 2 feet longer than tree height so (x + 2)² = 121 + x² x² + 4x + 4 = 121 +x² 4x = 117 x=29.25 (TREE HEIGHT) Length of Wire = 31.25 Double Check 31.25² = 29.25² + 11² 976.5625 = 855.5625 + 121 976.5625 = 976.5625